OpenStudy (anonymous):

simplify the expression √6 + 2√6

OpenStudy (anonymous):

I can take a crack at it if you like. Give me a minute! :)

OpenStudy (anonymous):

I believe you use the distributive property observed here: http://hotmath.com/tutor/?pid=genericalg1_8_1_RadicalExpressionsandRadicalEquations_15_700&gt_index=genericalg1:8:1:7

OpenStudy (anonymous):

Does this look familar?

OpenStudy (anonymous):

sorry! thats as much as i know :(

OpenStudy (anonymous):

its basically where you add the two numbers in front of the radical because they are like terms. if that makes sense. there is a understood 1 in front \[\sqrt{6} \]. its like when you add 2x and 4x. but you cannot add for example. \[2x ^{2} \] and 2x. These share the common term of radical 6.

OpenStudy (anonymous):

ok, what dont you understand? ill try my best to answer!

OpenStudy (anonymous):

Ok. you are using the distributive property. It looks like this: \[a \sqrt{b}+c \sqrt{b}=(a+c)\sqrt{b}\] Do you understand this?

OpenStudy (anonymous):

ok! you use it like this for this equation. \[1\sqrt{6}+2\sqrt{6}=(1+2)\sqrt{6}\]

OpenStudy (anonymous):

you had it right the first time! D:

OpenStudy (anonymous):

Nooo. i just meant that i couldnt explain WHY thats the answer.

OpenStudy (anonymous):

Im sorry. that is the correct answer.

OpenStudy (anonymous):

I hope this helped a little :( there is some great exercises online for these types of equations if you need it.